For a standard normal distribution, find cc if

P(z > c) = 0.3553.

How do you find this WITHOUT using the z table?
Answer #1

Solution :

P(z > c) = 0.3553

1 - P(z < c) = 0.3553

P(z < c) = 1 - 0.3553 = 0.6447

P(z < 0.3711) = 0.6447

c = 0.3711